Decompose the following into partial fractions

3)5x2+20x+6 / x3+2x2+x


4) x2+2x+3 / (x+1)(x2+2x+4)


1
Expert's answer
2021-09-06T07:23:00-0400

Solution 3:

"\frac{5 x^{2}+20 x+6}{x(x+1)^{2}}=\frac{A}{x}+\frac{B}{x+1}+\frac{C}{(x+1)^{2}}"  

"\\
\Rightarrow 5 x^{2}+20 x+6=A(x+1)^{2}+B x(x+1)+C x"  

Equating constants

"6=\mathrm{A}"  

Equating coefficients of  "\mathrm{x}^{2}"

"5=A+B\\

B=-1"

Equating coefficients of  "x"

"20=2 A+B+C\\
\Rightarrow 20=12-1+C \\
\Rightarrow C=9"  

"\therefore \frac{5 x^{2}+20 x+6}{x(x+1)^{2}}=\frac{6}{x}-\frac{1}{x+1}+\frac{9}{(x+1)^{2}}"



Solution 4:

"\frac{x^2+2x+3}{(x+1)(x^2+2x+4)}\\
=\frac{x^2+2x+4-1}{(x+1)(x^2+2x+4)}\\
=\frac{1}{x+1}-\frac{1}{(x+1)(x^2+2x+4)}\\"


Now decomposing "\frac{1}{(x+1)(x^2+2x+4)}\\" as: "\frac{A}{x+1}+\frac{Bx+C}{x^2+2x+4}"

"\therefore \frac{1}{(x+1)(x^2+2x+4)}=\frac{A}{x+1}+\frac{Bx+C}{x^2+2x+4}\\
\Rightarrow 1=A(x^2+2x+4)+(Bx+C)(x+1)\\
\Rightarrow 1=(A+B)x^2+(2A+B+C)x+4A+C"

"\therefore A+B=0;2A+B+C=0;4A+C=1\\
\Rightarrow A=-B \\
\therefore -2B+B+C=0\\
\Rightarrow B=C\\"

So, "4(-B)+B=1\\"

"\Rightarrow B=-\frac{1}{3}\\"

"\therefore A=\frac{1}{3},C=-\frac{1}{3}"

So, "\frac{1}{(x+1)(x^2+2x+4)}=\frac{1}{3(x+1)}-\frac{x+1}{3(x^2+2x+4)}"

Hence, "\frac{x^2+2x+3}{(x+1)(x^2+2x+4)}=\frac{4}{3(x+1)}-\frac{x+1}{3(x^2+2x+4)}"